A 50.0-g sample of copper is at 25.0°C. If 1 200 J of energy is added to it by heat, what is the final temperature of the copper?


1
Expert's answer
2021-05-17T05:50:02-0400

The amount of heat energy Q needed to raise the temperature of a substance is proportional to

the temperature change and to the mass of the substance:


"Q=m\cdot c \cdot \Delta T"

where c - the specific heat (the heat capacity per unit mass).

for Copper:


"c=0.386 \frac{J}{g \cdot K}=0.386 \frac{J}{g \cdot ^\circ C}"

then


"\Delta T=\frac{Q}{m \cdot c}=\frac{1200J}{50g \cdot 0.386 \frac{J}{g \cdot ^\circ C}} \approxeq 62.2 ^\circ C"

The final copper temperature will be


"T=25.0^\circ C+\Delta T=25.0^\circ C+62.2^\circ C=87.2^\circ C"

Answer: "87.2^\circ C"